OpenStudy (anonymous):

Solve the following system of equations. x + 4y + z = –10 3x – 3y + 6z = –21 x + 2y + 2z = –10 Answer (–4, –1, –2) (2, –2, –4) (–3, –2, –3) (1, –2, –3)

OpenStudy (anonymous):

Didn't you ask this already?

OpenStudy (anonymous):

yeah but i need help

OpenStudy (anonymous):

Yeah, but we were helping you and you left..

OpenStudy (anonymous):

So I'll ask again. Do you want help, or just the answer?

OpenStudy (anonymous):

And again you seem to have left...

OpenStudy (anonymous):

help to the answer so kinda both

OpenStudy (anonymous):

So what do you get if you subtract 3 times the 3rd equation from the second one?

OpenStudy (anonymous):

let me ask you this, what kind of math are you in david?

OpenStudy (anonymous):

x + 4y + z = –10 3x – 3y + 6z = –21 x + 2y + 2z = –10 (1 4 1 -10) (3 -3 6 -21) (1 2 2 -10) have you ever seen this type of notation david?

OpenStudy (anonymous):

algebra2

OpenStudy (anonymous):

ok do you see the first and third eq have the same answer -10. set those two equal to eachother and isolate one variable

OpenStudy (anonymous):

x+4y+z=x+2y+2z z=2y

OpenStudy (anonymous):

okey then what

OpenStudy (anonymous):

ok what you need to do is 1st pick two eq's solve for two variables and solve that. then pick two others and solve for one new variable and one old one. then sub those back in to solve the equations. its really hard to do it on this site. there should be an equation in algebra 2 that does this an easier way than elimination. but since they gave you the answers maybe they have not shown you how to solve for these things yet.